OpenStudy (anonymous):

Help Please! Reduce rational expression, ax to the 2 power b to the 2 power ____________________ . ______________________ b x to the 2 power y Thank you.

OpenStudy (anonymous):

ab/y

OpenStudy (saifoo.khan):

\[\frac{ax^2}{b} \times \frac{b^2}{x^2y}\]

OpenStudy (saifoo.khan):

\[\frac{ab}{y}\]

OpenStudy (anonymous):

ab over y is right.... 

OpenStudy (anonymous):

Thanks everyone. I thought it was ax over y... doesn't the 2 b's cross each other out?

OpenStudy (anonymous):

The two b's don't cross each other out because b^2 is a part of the numerator and b is a part of the denominator. If you cross out the b from the denom, and one b from the numerator, you still have one b left :)

OpenStudy (anonymous):

k, gotcha. thanks again :)

OpenStudy (anonymous):